OpenStudy (anonymous):

Solve the system of equations by the substitution method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. x + 3y = 32 -3x + 2y = 3 A. (4, 10) B. No solution; inconsistent C. (5, 9) D. (-5, 10)

OpenStudy (anonymous):

C. (5, 9) Would you like steps?

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

Stack the two equations. (1) x + 3y = 32 (2) -3x +2y = 3 Add the equations together as to cancel out one of the variables. You can multiply one of the equations to have them more easily cancel each other out. For example, we'll multiply equation 1 by 3 to yield: 3x + 9y = 96 Then add the two equations yielding: 11y = 99 y=9 Then plug and chug that into either equation and solve for x. Sometimes, you need to check if negative number work, so the answer could be more than one number.

OpenStudy (anonymous):

OH :o okay omg thank you so much :D!

OpenStudy (anonymous):

yw